OpenStudy (anonymous):

what percentage is the growth between 7100 and 8449?

OpenStudy (kropot72):

\[\large Percentage\ growth=\frac{8449-7100}{7100}\times\frac{100}{1}=you\ can\ calculate\]

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome :)